OpenStudy (anonymous):

Box A has a volume of 12 cubic meters. Box B is similar to box A. To create box B, box A's dimensions were multiplied by five. What is the volume of box B? a. 60 m3 b. 300 m3 c. 1,500 m3 d. 7,500 m3

OpenStudy (anonymous):

We're going to assume that box A is a cube to make things easy. You get the volume of a cube, or any box shaped thing, by multiplying the width by the length by the height. So, you know that Volume = W*L*H and since this is a cube, H = W = L, you get the equation Volume = W^3 (you could use L or H too), so, to find W, you need to cube root Volume. \[\sqrt[3]{12} = W\] Since box A and B are similar, we can just multiply W by 5 and then cube it to get the answer. Should be C.